OpenStudy (anonymous):

Newton's Law of Gravitation. Given that G= 6.674 10^ -11N m^2/kg^2, ME= 5.98 10^24 kg, and RE= 6380 km… Calculate a theoretical value for g.

OpenStudy (anonymous):

\[F = ma \] \[F_g = m_{you!} a_g\] \[F_g = G\frac{m_1m_2}{r^2}= G\frac{ \ m_{you!} \ M_E}{R_E^2}\]\] \[G\frac{\cancel{m}M_E}{R_E^2} = \cancel{m}a_g\] \[a_g = G\frac{M_E}{R_E^2}\]